what is the product of 5 and some number decreased by 7 is equal to 12?
Answer by swincher4391(1107)

Some number (n) times 5 plus 7 gives us 12. This is simply:
5n + 7 = 12
and if you want to solve:
5n +7-7 = 12 -7
5n = 5
n = 1